1. Market equilibrium is one of the fundamental concepts in economics.
  2. Describe market equilibrium using relevant graphs and discuss market surplus and market shortage. 
  3. Suppose the market for kittens from an adoption centre can be described by the following equations:

 

Demand equation: Qd = 100 – 20P

Supply equation: Qs = 130 + 2P

 

Calculate the equilibrium price (P) and quantity (Q) of kittens. Remember that a negative price for kittens is not allowed. How many kittens will be adopted by humans and how many will be “strays?” 






1
Expert's answer
2021-05-25T16:57:58-0400

Market Equilibrium

This is a situation in the economy whereby market forces are balanced such that quantity demanded is equal to quantity supplied.

For example in the graph the quantity of gasoline supplied and quantity demanded are equal where the curves intersect

Market Surplus

Market Surplus occurs when the quantity supplied is more than the quantity demanded. This happens when the price of the commodity is higher than the equilibrium price.

An example is shown in the figure below


Market Shortage

This happens when the quantity supplied into the market is less than the quantity demanded. This normally happens when the price of the commodity is less than the equilibrium price. Therefore manufacturers opt to supply less whereas consumers demand more. The deficit between the quantity demanded and quantity supplied is the Market Shortage.

As shown below.


3.Demand equation: Qd = 100 – 20P

Supply equation: Qs = 130 + 2P

Calculating equilibrium price and quantity of kittens

Qs=Qd

"100-20P= 130+2P"

"100-130= 2P+20P"

"-30=22P"

"P=\frac{30}{22}"

"P= 1.36"


Quantity at equilibrium

We replace the value of P in any of the two equation to obtain quantity at equilibrium

Thus,

"100-20(-1.36)"

"Quantity =127"

Note

The negative sign is used while calculating quantity but ignored when quoting the price( there is no negative cost).